Q.1) Fresh Food has recently carried out a survey of the demand for their seasonal fruits. They observed the following

outcomes over the last seven months:

  Quantity Sold

  8500

  4700

  5800

  7400

  6200

  7300

  5600

Price Advertising

$ 2 $ 2,800 $ 5 $ 200 $ 3 $ 400 $ 2 $ 500 $ 5 $ 3,200 $ 3 $ 1,800 $ 4 $ 900

            Regression (Equation)

QSold = 8536.213882 - 835.7223514 Price + 0.592228496 Advertising

A. Estimate sales when the price is $50 and advertising is $100,000, stating any assumption that you need to make.

B. If the firm charges $50 and cost per unit is $30 but increases advertising to $110,000, what conclusions can you derive in terms of revenues and profits?

Q


1
Expert's answer
2022-03-29T12:13:07-0400

A: "QSold = 8536.213882 - 835.7223514 Price + 0.592228496 Advertising"

price = $50, Advertising is =$100,000

"QSold = 8536.213882 - 835.7223514 \times50 + 0.592228496\times100000"

"Qsold= 25,972.945912"


B: Quantity sold at $50 if advertising is 110000

"QSold = 8536.213882 - 835.7223514 \times50 + 0.592228496\times110000\\
Qsold= 31,895.230872\\total\:sales=31,895.230872\times50=1,594,761.5436\\cost \:of\:goods=31,859.230872\times30=956,856.92616"

The profits and revenue increase because of increased sales arising from increased advertising.